OpenStudy (anonymous):

Which of the following describes acceleration? 1:s/m 2:m/\(\sf s^2\) 3:\(\sf m^2\) 4:m/s

OpenStudy (abhisar):

Hi! Welcome to OpenStudy! What do you think it should be?

OpenStudy (abhisar):

Reload the page to see the question correctly.

jhonyy9 (jhonyy9):

from physic knowledge i think 2. is right

OpenStudy (marihelenh):

That is correct because acceleration is equal to velocity over time. The units for velocity are meters per second and time is seconds, so you have, \[\frac{ \frac{ m }{ s } }{ s }\] Instead of dividing, you can multiply by switching the numerator and denominator, so \[\frac{ m }{ s }*\frac{ 1 }{ s }\]

OpenStudy (marihelenh):

That will give you your units, which the second option is the only one with the correct units.